OpenStudy (anonymous):

Find values of a & b that make each set of equations true. 6=a•b^0. 32=a•b^2 24=a•b^2. 128=a•b^3

OpenStudy (anonymous):

b^0 = 1 Therefore, a = 6 Substitute a = 6 into 24 = a x b^2 24 = 6 x b^2 24/6 = b^2 4 = b^2 sqrt(4) = b 2 = b BUT why is 32 = a x b^2 the same as 24 = a x b^3 with the expression, are you sure that you haven't mis-typed?